The utility function of the consumer is 5x
0.5y
0.5

The income of the consumer is 5000
The price of good x is 1,000 and the price of good y is 500.
Determine the optimal consumption bundle.
1
Expert's answer
2020-10-20T06:28:50-0400

Solution:

Utility Function:"U_(x,y) =5x^{0.5} y^{0.5}"


"MRS_{x,y} =\frac{MUx}{MUy} \;where\;MUx=\frac{dU}{dy} =2.5x^{-0.5} y^{0.5}"


"where\;MUy=\frac{dU}{dy} =0.5x^{0.5} y^{-0.5}"


"MRS_{x,y} =\frac{MUx}{MUy}"


"=\frac{2.5x^{-0.5} y^{0.5} }{0.5x^{0.5} y^{-0.5} } =\frac{2.5x(\frac{y}{x} )^{0.5} }{0.5x(\frac{x}{y}) ^{0.5}}=5(\frac{y}{x})"


"MRS=5(\frac{y}{x})"


"We \;know\; that\; MRS=\frac{Px}{Py}"


"5(\frac{y}{x})=\frac{1000}{500}"


"5(\frac{y}{x})=2"


"y=\frac{2}{5} x"


Plug this into the budget line:

Budget Line:


"I=PxX+PyY"

"5000=1000X+500Y"

"5000=1000X+500(\frac{2}{5} x)"


"5000=1000X+200X"

"5000=1200X"

"X=\frac{5000}{1200}"


"X=4.17"


Plug X into the budget line to get the value of Y:


"5000=1000X+500Y"

"5000=1000(4.17) +500Y"

"5000=4170 +500Y"

"5000-4170=500Y"

"830=500Y"


"Y=\frac{830}{500}"


"Y=1.66"


The optimum consumption bundle is therefore (x,y) = (4.17, 1.66)